Can someone please help! I need to solve for 'w' but i dont know how to solve it, can someone show me the steps on how to solve.
(2/8)+(3/4)= (w/6)

Answer by rfer(16322)

2/8+3/4=w/6
2/8+6/8=w/6
8/8=w/6
8/8=1
1=w/6
1*6=w
6=w